OpenStudy (anonymous):

find the integral of xe^x from -infinity to 0

OpenStudy (anonymous):

\[\int\limits_{-\infty}^{0}xe ^{x}dx\]

OpenStudy (helder_edwin):

\[ \large \int_{-\infty}^0xe^x\,dx=\lim_{b\to+\infty}\int_{-b}^0xe^x\,dx \]

OpenStudy (helder_edwin):

use integration by parts

OpenStudy (anonymous):

I used integration by parts and got stuck at a particular place towards the end.

OpenStudy (anonymous):

u = x,u'=1,v=e^x,v'=e^x

OpenStudy (helder_edwin):

\[ \large \begin{alignat*}{2} u&=x &\qquad dv=e^x\\ du&=dx&\qquad v=e^x \end{alignat*} \] so \[ \large \int xe^x\,dx=xe^x-\int e^x\,dx=xe^x-e^x \]

OpenStudy (anonymous):

ok so then would you take the limit of xe^x - e^x?

OpenStudy (helder_edwin):

right \[ \large \lim_{b\to+\infty}\int_{-b}^0xe^x\,dx= \lim_{b\to+\infty}\left[ xe^x-e^x\right]_b^0 \]

OpenStudy (anonymous):

The main thing I don't understand is when you are solving with integration by parts is why you son't leave the domain for the integrals from -infinity to 0

OpenStudy (helder_edwin):

just to save time when typing

OpenStudy (anonymous):

because I would think you then would have to evaluate e^x from -infinity to 0 before you could take the limit. of xe^x - e^x

OpenStudy (anonymous):

ok but uv' evaluated at -inf to 0 = xe^x - [e^x evaluated at -inf to 0] ?

OpenStudy (helder_edwin):

\[ \large \int_{-\infty}^0xe^x\,dx=xe^x|_{-\infty}^0-\int_{-\infty}^0e^x\,dx \] i don't see where is the problem.

OpenStudy (helder_edwin):

\[ \large =xe^x|_{-\infty}^0-e^x|_{-\infty}^0=[xe^x-e^x]_{-\infty}^0 \]

OpenStudy (anonymous):

ah I think I understand now. thinking into it more. I appreciate your help.

OpenStudy (helder_edwin):

u r welcome